OpenStudy (anonymous):

Find the product of (x − 5)2.

OpenStudy (campbell_st):

well if its \[(x -5)^2 = (x -5) \times (x -5)\] a simple way to do this is to split the 1st binomial \[(x-5)^2 = x(x-5) -5(x -5)\] distribute and collect like terms for the solution

OpenStudy (anonymous):

thank you :)